### v4

# Algorithms for Approximating Conditionally Optimal Bounds

Algorithms for Approximating Conditionally Optimal Bounds
George Bissias
(March 2025)
1 Preliminaries
In this document we consider lower bounds on discrete distributions $\mathcal{F}^{m}$ over support set
$S^{m}=\left\{S_{\min}+i\frac{S_{\max}-S_{\min}}{m-1}~\bigg|~i\in\mathbb{N},0\leq i\leq m-1\right\}.$ (1)
Samples of size $n$ are drawn from sample space $\Omega^{m}=S^{m}\times\ldots\times S^{m}$. We further assume that samples are unique when their components are arranged in increasing order. When a statement is true for any $m$, we often drop the superscript from the associated quantities.
Our results concern bounds consistent with preorders on the samples in $\Omega$. A preorder can be characterized by a set of equivalence classes, combined with a partial order over those classes. In the case where every equivalence class contains a single sample, and the classes are totally ordered, the preorder is also a total order. Learned-Miller [2] formalized the notion of consistency for total orders and we extend that theory to preorders.
Definition 1.
For any given preorder $R$ and samples ${\bm{x}},{\bm{y}}\in\Omega$, we write ${\bm{x}}\lesssim_{R}{\bm{y}}$ whenever $\bm{y}$} is greater than or equal to $\bm{x}$} in $R$. A preorder is characterized by the following properties
  • ${\bm{x}}\lesssim_{R}{\bm{x}}$.
  • ${\bm{x}}\lesssim_{R}{\bm{y}}$ and ${\bm{y}}\lesssim_{R}\bm{z}$ implies that ${\bm{x}}\lesssim_{R}\bm{z}$.
When ${\bm{x}}\lesssim_{R}{\bm{y}}$ and ${\bm{y}}\lesssim_{R}{\bm{x}}$ we write ${\bm{x}}\sim_{R}{\bm{y}}$. And by ${\bm{x}}<_{R}{\bm{y}}$ we indicate that ${\bm{x}}\lesssim_{R}{\bm{y}}$ and ${\bm{x}}\not\sim_{R}{\bm{y}}$.
Definition 2.
Partial order $R$ is a preorder where $\forall{\bm{x}},{\bm{y}}\in\Omega,{\bm{x}}\sim_{R}{\bm{y}}$ implies that ${\bm{x}}={\bm{y}}$. When $\bm{x}$} is no greater than $\bm{y}$} in partial order $R$ we write ${\bm{x}}\leq_{R}{\bm{y}}$. We further user ${\bm{x}}<_{R}{\bm{y}}$ to indicate that ${\bm{x}}\leq_{R}{\bm{y}}$ and ${\bm{x}}\neq{\bm{y}}$.
Definition 3.
Total order $T$ is a partial order that is strongly connected, i.e., $\forall{\bm{x}},{\bm{y}}\in\Omega,{\bm{x}}\leq_{T}{\bm{y}}$ or ${\bm{y}}\leq_{T}{\bm{x}}$.
Definition 4.
Total preorder $R$ is a preorder that is strongly connected, i.e. where $\forall{\bm{x}},{\bm{y}}\in\Omega,{\bm{x}}\lesssim_{R}{\bm{y}}$ or ${\bm{y}}\lesssim_{R}{\bm{x}}$. A total preorder is a total order over equivalence classes.
Proposition 1.
If $R$ is a preorder, then for any ${\bm{x}},{\bm{y}}\in\Omega$. We have
${\bm{x}}<_{R}{\bm{y}}\Rightarrow\neg({\bm{y}}\lesssim_{R}{\bm{x}}).$ (2)
If additionally $R$ is a total preorder, then we also have
$\neg({\bm{y}}\lesssim_{R}{\bm{x}})\Rightarrow{\bm{x}}<_{R}{\bm{y}}.$ (3)
Proof.
To prove the forward direction, suppose for the sake of contradiction that ${\bm{x}}<_{R}{\bm{y}}$ and ${\bm{y}}\lesssim_{R}{\bm{x}}$. Since ${\bm{x}}<_{R}{\bm{y}}$ we know that ${\bm{x}}\lesssim_{R}{\bm{y}}$ and ${\bm{x}}\not\sim_{R}{\bm{y}}$. But since ${\bm{x}}\lesssim_{R}{\bm{y}}$ and ${\bm{y}}\lesssim_{R}{\bm{x}}$ we have by Definition 1 that ${\bm{x}}\sim_{R}{\bm{y}}$, which contradicts the earlier observation that ${\bm{x}}\not\sim_{R}{\bm{y}}$.
To prove the reverse direction, suppose that $R$ is a total preorder and $\neg({\bm{y}}\lesssim_{R}{\bm{x}})$. Since $R$ is a total preorder, we know by Definition 4 that either ${\bm{x}}\lesssim_{R}{\bm{y}}$ or ${\bm{y}}\lesssim_{R}{\bm{x}}$. Therefore, it must be the case that ${\bm{x}}\lesssim_{R}{\bm{y}}$. Now Definition 1 requires that ${\bm{x}}\lesssim_{R}{\bm{y}}$ and ${\bm{y}}\lesssim_{T}{\bm{x}}$ in order for ${\bm{y}}\sim_{R}{\bm{x}}$. Therefore ${\bm{x}}\not\sim_{R}{\bm{y}}$, since $\neg({\bm{y}}\lesssim_{R}{\bm{x}})$. Thus, ${\bm{x}}\lesssim_{R}{\bm{y}}$ and ${\bm{x}}\not\sim_{R}{\bm{y}}$ so that ${\bm{x}}<_{R}{\bm{y}}$ by Definition 1. ∎
Definition 5.
The upper set associated with sample $\bm{x}\in\Omega$ and total preorder $R$ is given by
$\Omega(\bm{x},R)=\{{\bm{y}}\in\Omega:\bm{x}\lesssim_{R}\bm{y}\}.$ (4)
Note that, according to these definitions, $\forall\bm{x},\bm{y}\in\Omega$, total preorder $R$, and total order $T$, $\bm{x}\lesssim_{R}\bm{y}\Leftrightarrow\bm{y}\in\Omega(\bm{x},R)$ and ${\bm{x}}\leq_{T}\bm{y}\Leftrightarrow\bm{y}\in\Omega(\bm{x},T)$.
Definition 6.
For $\bm{x},\bm{y}\in\Omega$ we write $\bm{x}\leq\bm{y}$ whenever $\forall i\in[n],x_{(i)}\leq y_{(i)}$. We write $\bm{x}<\bm{y}$ whenever $\bm{x}\neq\bm{y}\wedge\bm{x}\leq\bm{y}$.
Definition 6 implies that $\neg(\bm{x}\leq\bm{y})\Leftrightarrow\exists i\in[n]:x_{(i)}>y_{(i)}$, and $\neg(\bm{x}<\bm{y})\Leftrightarrow(\bm{x}=\bm{y})\vee\neg(\bm{x}\leq\bm{y})$. Furthermore,
$\bm{x}<\bm{y}\Rightarrow\exists i\in[n]:x_{(i)}<y_{(i)}\Rightarrow\neg(\bm{y}\leq\bm{x}).$ (5)
Definition 7.
By $\bm{S}_{i}^{m}$ we denote the homogeneous sample of length $n$:
$(S_{i}^{m},\ldots,S_{i}^{m})\in\Omega^{m}$ for $S_{i}^{m}\in S^{m}$. We will frequently drop the superscript $m$ from $\bm{S}^{m}_{i}$ when it appears in a statement that is true for any $m$.
1.1 Common orders
Definition 8.
The low lexicographic order, $T_{\ell}$, is the total order such that $\forall\bm{x},\bm{y}\in\Omega,\bm{x}\leq_{T_{\ell}}\bm{y}\Leftrightarrow L_{1}(\bm{x},\bm{y})=1$ where $\forall i\in[n]$,
$L_{i}(\bm{x},\bm{y})=(x_{(i)}<y_{(i)})\vee(x_{(i)}=y_{(i)})\wedge L_{i+1}(\bm{x},\bm{y}),$ (6)
and $L_{n+1}(\bm{x},\bm{y})=1$.
Definition 9.
The high lexicographic order, $T_{h}$, is the total order such that $\forall\bm{x},\bm{y}\in\Omega,\bm{x}\leq_{T_{h}}\bm{y}\Leftrightarrow H_{n}(\bm{x},\bm{y})=1$ where $\forall i\in[n]$,
$H_{i}(\bm{x},\bm{y})=(x_{(i)}<y_{(i)})\vee(x_{(i)}=y_{(i)})\wedge H_{i-1}(\bm{x},\bm{y}),$ (7)
and $H_{0}(\bm{x},\bm{y})=1$.
Definition 10.
For fixed $i\in\{1,\ldots,n\}$ and any two samples ${\bm{x}},{\bm{y}}\in\Omega$, the $i$th quantile preorder $R_{i}$ is the total preorder defined by
${\bm{x}}\lesssim_{R_{i}}{\bm{y}}\Leftrightarrow x_{(i)}\leq y_{(i)}.$ (8)
Notice that when $n$ is odd and $i=\lceil n/2\rceil$, the $i$th quantile is equal to the sample median.
Definition 11.
Preorder $R$ is weakly monotone if $\bm{x}\leq\bm{y}\Rightarrow\bm{x}\lesssim_{R}\bm{y}$, and it is strongly monotone if $\bm{x}<\bm{y}\Rightarrow\bm{x}<_{R}\bm{y}$.
Proposition 2.
Every strongly monotone order is also weakly monotone.
Proof.
Suppose that $R$ is a strongly monotone order and take any ${\bm{x}},{\bm{y}}\in\Omega$ such that ${\bm{x}}\leq{\bm{y}}$. In the case that ${\bm{x}}<{\bm{y}}$ we know that ${\bm{x}}<_{R}{\bm{y}}$ as well by the strong monotonicity property, which implies that ${\bm{x}}\lesssim_{R}{\bm{y}}$. Now suppose that ${\bm{x}}={\bm{y}}$. Then by the reflexive property of preorders, we have that ${\bm{x}}\sim_{R}{\bm{y}}$, so that ${\bm{x}}\lesssim_{R}{\bm{y}}$. In either case, ${\bm{x}}\leq{\bm{y}}\Rightarrow{\bm{x}}\lesssim_{R}{\bm{y}}$. ∎
Proposition 3.
A weakly monotone partial order is also strongly monotone.
Proof.
Let $R$ be an arbitrary weakly monotone partial order. By Definitions 2 and 6 we have that
$\bm{x}<\bm{y}\Rightarrow({\bm{x}}\leq\bm{y})\wedge({\bm{x}}\neq{\bm{y}})\Rightarrow({\bm{x}}\leq_{R}\bm{y})\wedge({\bm{x}}\neq{\bm{y}})\Rightarrow({\bm{x}}<_{R}\bm{y}).$ (9)
Proposition 4.
A weakly monotone preorder is strongly monotone iff $\forall{\bm{x}},{\bm{y}}\in\Omega,{\bm{x}}<{\bm{y}}\Rightarrow{\bm{x}}\not\sim_{R}{\bm{y}}$.
Proof.
Let $R$ be an arbitrary weakly monotone preorder. It’s clear that the conclusion is necessary for strong monotonicity. For suppose that it is false, i.e. there exists ${\bm{x}},{\bm{y}}\in\Omega$ where ${\bm{x}}<{\bm{y}}$ and ${\bm{x}}\sim_{R}{\bm{y}}$, then $R$ cannot be strongly monotone since we could not have ${\bm{x}}<_{R}{\bm{y}}$.
To show that the conclusion is sufficient for strong monotonicity notice that, since $R$ is weakly monotone by assumption, we have that
$\bm{x}<\bm{y}\Rightarrow{\bm{x}}\leq\bm{y}\Rightarrow{\bm{x}}\lesssim_{R}\bm{y}.$ (10)
And also by assumption, ${\bm{x}}\not\sim_{R}{\bm{y}}$, which implies that ${\bm{x}}<_{R}{\bm{y}}$. Thus, ${\bm{x}}<{\bm{y}}\Rightarrow{\bm{x}}<_{R}{\bm{y}}$. ∎
All monotone orders discussed in this document are strongly monotone. For brevity, we refer to such orders as simply monotone going forward.
2 Topology
Our results rely on the topological characteristics of the space of probability distributions supported on $S$. Because we deal exclusively with finite support, we may characterize the set of all distributions having support $S$ by the simplex of vectors $\Delta$, which are those ordered vectors in $0,1]^{m}$ having unit L1 norm; for ${\bm{u}}\in\Delta$, $u_{i}$ dictates the probability mass assigned to $S_{i}$. The natural topology on this space is the one induced by the Euclidean distance metric: $\forall{\bm{u}},{\bm{v}}\in\Delta,d({\bm{u}},{\bm{v}})=\|{\bm{u}}-{\bm{v}}\|_{2}$.
Definition 12.
For each $\alpha\in[0,1)$ and upper set $\Omega^{\prime}\subseteq\Omega$, the interior likely set is given by
$\mathcal{G}(\Omega^{\prime},\alpha)=\{F\in\mathcal{F}:P_{F}[\Omega^{\prime}]>\alpha\},$ (11)
and the likely set is given by
$\mathcal{F}(\Omega^{\prime},\alpha)=\texttt{cl}(\mathcal{G}(\Omega^{\prime},\alpha)),$ (12)
where cl denotes the closure of a set.
Let $\bm{s}$ denote the ordered vector of support points $S$ and note that for each ${\bm{u}}\in\Delta$, corresponding to distribution $H_{{\bm{u}}}$, $E[H_{{\bm{u}}}]={\bm{u}}\cdot\bm{s}$. Thus, for ${\bm{u}},{\bm{v}}\in\Delta$, corresponding to distributions $H_{{\bm{u}}}$ and $H_{{\bm{v}}}$ and such that $d({\bm{u}},{\bm{v}})<\delta$, we have that
$\begin{array}[]{rcl}|E[H_{{\bm{u}}}]-E[H_{{\bm{v}}}]|&\leq&S_{\max}\left|\sum_{i=1}^{m}u_{i}-v_{i}\right|\\ &\leq&S_{\max}\|{\bm{u}}-{\bm{v}}\|_{1}\\ &\leq&S_{\max}\sqrt{m}\|{\bm{u}}-{\bm{v}}\|_{2}\\ &=&d({\bm{u}},{\bm{v}})\sqrt{m}S_{\max}\\ &<&\delta\sqrt{m}S_{\max}\\ \end{array}.$ (13)
Lemma 1.
Fix $\alpha\in[0,1)$, $\Omega^{\prime}\subseteq\Omega$, and $\delta>0$. For each $F\in\mathcal{F}(\Omega^{\prime},\alpha)$ there exists a $G\in\mathcal{G}(\Omega^{\prime},\alpha)$ such that $|E[F]-E[G]|<\epsilon$.
Proof.
Take any $F\in\mathcal{F}(\Omega^{\prime},\alpha)$. By the definition of closure we know that $\forall\delta>0,\exists G\in\mathcal{G}(\Omega^{\prime},\alpha):d(G,F)<\delta$. Choose $\delta=\epsilon/(\sqrt{m}S_{\max})$. According to Inequality 13, we have $|E[F]-E[G]|<\epsilon$. ∎
Lemma 2.
Take any $\alpha\in[0,1)$, $\mu\in[S_{\min},S_{\max}]$, and $\Omega^{\prime}\subseteq\Omega$. Then there exists $F\in\mathcal{F}(\Omega^{\prime},\alpha)$ such that $E[F]=\mu$ if and only if there exists a sequence of distributions in $\mathcal{G}(\Omega^{\prime},\alpha)$ that converge in mean to $\mu$.
Proof.
Suppose first that there exists $F\in\mathcal{F}(\Omega^{\prime},\alpha)$ such that $E[F]=\mu$ and choose any $\epsilon>0$. By the definition of closure we know that $\forall i>0,\exists G_{i}\in\mathcal{G}(\Omega^{\prime},\alpha):d(G_{i},F)<1/i$. According to Inequality 13, when $i>\sqrt{m}S_{\max}/\epsilon$, $|E[F]-E[G_{i}]|<\epsilon$. Thus, allowing $\epsilon$ to approach 0, we can see that the sequence of distributions $\{G_{i}\}_{i>0}$ converges in mean to $F$.
Next, suppose that there exists a sequence of distributions $\{G_{i}\}_{i>0},G_{i}\in\mathcal{G}(\Omega^{\prime},\alpha)$ converging in mean to $\mu$. It suffices for the limit of this sequence to be in $\mathcal{F}(\Omega^{\prime},\alpha)$, which is true because $\lim_{i\rightarrow\infty}\{G_{i}\}_{i>0}$ is a limit point of $\mathcal{G}(\Omega^{\prime},\alpha)$, and the closure of any set contains all of its limit points. ∎
3 Conditional-Optimality
Definition 13.
Preorder order $D$ agrees with a preorder $R$ if $\forall{\bm{x}},{\bm{y}}\in\Omega,{\bm{x}}<_{R}{\bm{y}}\Rightarrow{\bm{x}}<_{D}{\bm{y}}$. Note that agreement is not necessarily a symmetric property.
Proposition 5.
For every preorder $R$, there exists a total order $T$ that agrees with $R$.
Proof.
We proceed in two stages by first showing that there exists a partial order that agrees with $R$ and then invoking the Szpilrajn extension theorem [3] to argue that there exists a total order that agrees with the partial order.
Every preorder $R$ is a partial order over equivalence classes. Construct partial order $D$ by maintaining all non-equivalence relations of $R$ and arbitrarily assigning a total order between the elements within each of the equivalence classes of $R$ (that such a total order exists follows from the well-ordering principle). Now take any ${\bm{x}},{\bm{y}}\in\Omega$ such that ${\bm{x}}<_{R}{\bm{y}}$. Since ${\bm{x}}\not\sim_{R}{\bm{y}}$, it’s clear by construction that ${\bm{x}}<_{D}{\bm{y}}$ as well. Thus, partial order $D$ agrees with $R$.
The Szpilrajn extension theorem [3] establishes that there must also exist a total order $T$ such that $\forall{\bm{x}},{\bm{y}}\in\Omega,{\bm{x}}\leq_{D}{\bm{y}}\Rightarrow{\bm{x}}\leq_{T}{\bm{y}}$. Fix ${\bm{x}},{\bm{y}}\in\Omega$ such that ${\bm{x}}<_{D}{\bm{y}}$ and let $T$ be the total order guaranteed by the theorem. We must then have that ${\bm{x}}\leq_{T}{\bm{y}}$. And of course ${\bm{x}}\neq{\bm{y}}$ by Definition 2. Thus, by Definition 3, ${\bm{x}}<_{T}{\bm{y}}$ as well. Therefore, total order $T$ agrees with $D$. ∎
Proposition 6.
If total order $T$ agrees with total preorder $R$, then $\forall{\bm{x}}\in\Omega$, $\Omega({\bm{x}},T)\subseteq\Omega({\bm{x}},R)$.
Proof.
Let ${\bm{x}}\in\Omega$ be arbitrary and note that by Definition 13, if total order $T$ agrees with total preorder $R$, then for each ${\bm{y}}\in\Omega$
$\neg({\bm{y}}<_{T}{\bm{x}})\Rightarrow\neg({\bm{y}}<_{R}{\bm{x}}).$ (14)
It follows that
[…]
Proposition 7.
For every total order $T$ agreeing with total preorder $R$ and any ${\bm{x}}\in\Omega$, there exists a ${\bm{y}}\in\Omega$ such that ${\bm{x}}\sim_{R}{\bm{y}}$ and $\Omega({\bm{x}},R)=\Omega({\bm{y}},T)$.
Proof.
Fix ${\bm{x}}\in\Omega$ and take $\bm{y}$} such that ${\bm{x}}\sim_{R}{\bm{y}}$ and $\forall{\bm{z}}\in\Omega:{\bm{x}}\sim_{R}{\bm{z}}$, ${\bm{y}}\leq_{T}{\bm{z}}$, i.e. $\bm{y}$} is the smallest element according to $T$ that is equivalent to $\bm{x}$} according to $R$. Such a $\bm{y}$} must exist by the well-ordering principle. By Definition 5, $\Omega({\bm{x}},R)=\Omega({\bm{y}},R)$ so that, invoking Proposition 6, it suffices to show that $\Omega({\bm{y}},R)\subseteq\Omega({\bm{y}},T)$.
We can partition $\Omega({\bm{y}},R)$ such that $\Omega({\bm{y}},R)=S_{1}\cup S_{2}$ where $S_{1}=\{{\bm{z}}:{\bm{y}}\sim_{R}{\bm{z}}\}$ and $S_{2}=\{{\bm{z}}:{\bm{y}}<_{R}{\bm{z}}\}$. Now take any ${\bm{z}}\in\Omega({\bm{y}},R)$. If ${\bm{z}}={\bm{y}}$, then it’s clear by Definition 5 that ${\bm{z}}\in\Omega({\bm{y}},T)$. If ${\bm{z}}\in S_{1}\setminus\{{\bm{y}}\}$, then ${\bm{y}}<_{T}{\bm{z}}$ by our choice of $\bm{y}$}, so that, again by Definition 5, ${\bm{z}}\in\Omega({\bm{y}},T)$. Finally, suppose that ${\bm{z}}\in S_{2}$. Since $T$ agress with $R$, we have by Definition 13 that ${\bm{y}}<_{T}{\bm{z}}$. Thus, by Definition 5, ${\bm{z}}\in\Omega({\bm{y}},T)$, and therefore $\Omega({\bm{y}},R)\subseteq\Omega({\bm{y}},T)$. ∎
Definition 14.
A bound $B$ is consistent with total order $T$ if $\forall{\bm{x}},{\bm{y}}\in\Omega$,
${\bm{x}}\leq_{T}{\bm{y}}\Rightarrow B({\bm{x}})\leq B({\bm{y}}).$ (16)
Definition 15.
Bound $B$ is consistent with preorder $R$ if it is consistent with every total order that agrees with $R$.
Lemma 3.
If bound $B$ is consistent with preorder $R$ then for any ${\bm{x}},{\bm{y}}\in\Omega$, $B({\bm{x}})\leq B({\bm{y}})$ whenever ${\bm{x}}<_{R}{\bm{y}}$ and $B({\bm{x}})=B({\bm{y}})$ whenever ${\bm{x}}\sim_{R}{\bm{y}}$. The converse also holds if $R$ is a total preorder.
Proof.
Suppose $B$ is consistent with preorder $R$ and take first any ${\bm{x}},{\bm{y}}\in\Omega$ where ${\bm{x}}<_{R}{\bm{y}}$. Definition 15 dictates that $B$ must be consistent with every total order $T$ agreeing with $R$. And Definition 13 requires that we must also have ${\bm{x}}<_{T}{\bm{y}}$, which in turn implies, by Definition 14, that $B({\bm{x}})\leq B({\bm{y}})$. Next, take ${\bm{x}},{\bm{y}}\in\Omega$ such that ${\bm{x}}\sim_{R}{\bm{y}}$. Let $T$ be any total order agreeing with $R$ such that ${\bm{x}}<_{T}{\bm{y}}$, and let $T^{\prime}$ be the total order identical to $T$ except that it reverses the order of $\bm{x}$} and $\bm{y}$}. Since $B$ is consistent with $R$, Definition 15 ensures that $B$ is consistent with both $T$ and $T^{\prime}$. This implies that $B({\bm{x}})\leq B({\bm{y}})$ and $B({\bm{y}})\leq B({\bm{x}})$ so that $B({\bm{x}})=B({\bm{y}})$.
Now suppose that $R$ is a total preorder, $B({\bm{x}})=B({\bm{y}})$ whenever ${\bm{x}}\sim_{R}{\bm{y}}$, and $B({\bm{x}})\leq B({\bm{y}})$ whenever ${\bm{x}}<_{R}{\bm{y}}$ for preorder $R$. Let $T$ be any total order agreeing with $R$. For each ${\bm{x}},{\bm{y}}\in\Omega$ we have by Proposition 1 and taking the contrapositive of Definition 13 that
${\bm{x}}\leq_{T}{\bm{y}}\Rightarrow\neg({\bm{y}}<_{T}{\bm{x}})\Rightarrow\neg({\bm{y}}<_{R}{\bm{x}})\Rightarrow{\bm{x}}\lesssim_{R}{\bm{y}}.$ (17)
If ${\bm{x}}<_{R}{\bm{y}}$, then by assumption $B({\bm{x}})\leq B({\bm{y}})$. And if ${\bm{x}}\sim_{R}{\bm{y}}$, we again have by assumption that $B({\bm{x}})=B({\bm{y}})\Rightarrow B({\bm{x}})\leq B({\bm{y}})$. Thus, in any case, ${\bm{x}}\lesssim_{R}{\bm{y}}\Rightarrow B({\bm{x}})\leq B({\bm{y}})$. It follows then by Implication 17 that ${\bm{x}}\leq_{T}{\bm{y}}\Rightarrow B({\bm{x}})\leq B({\bm{y}})$. Therefore, by Definition 14, it must be the case that $B$ is consistent with $T$. Since $T$ was an arbitrary total order agreeing with $R$, we have, by Definition 15, that $B$ is also consistent with $R$. ∎
Definition 16.
Bound $B$ consistent with preorder $R$ is conditionally optimal with respect to $R$ if for any other bound $B^{\prime}$ also consistent with $R$ we have $\forall{\bm{x}}\in\Omega,B^{\prime}({\bm{x}})\leq B({\bm{x}})$, and $\exists{\bm{y}}\in\Omega:B^{\prime}({\bm{y}})<B({\bm{y}})$.
Definition 17.
Let ${\bm{x}}\in\Omega$ and $\alpha\in[0,1)$ be arbitrary. The pessimal bound with respect to preorder $R$ is given by
$B_{R}^{*}({\bm{x}})=\min\{E[F]:F\in\mathcal{F}(\Omega(\bm{x},R),\alpha)\}.$ (18)
Theorem 1 (Learned-Miller [2]).
Let ${\bm{x}}\in\Omega$, $\alpha\in[0,1)$, and total order $T$ be arbitrary. When $\mathcal{F}(\Omega(\bm{x},T),\alpha)$ is non-empty, the pessimal bound $B_{T}^{*}({\bm{x}})$ is conditionally optimal with respect to $T$.
For the remainder of this document we will assume that for any preorder $R$ under consideration, $\alpha$ is chosen so that $\mathcal{F}(\Omega(\bm{x},R),\alpha)$ is indeed non-empty.
Lemma 4.
For any total preorder $R$, $B_{R}^{*}$ is consistent with $R$.
Proof.
According to Lemma 3, it will suffice to show that $B_{R}^{*}({\bm{x}})\leq B_{R}^{*}({\bm{y}})$ whenever ${\bm{x}}<_{R}{\bm{y}}$ and $B_{R}^{*}({\bm{x}})=B_{R}^{*}({\bm{y}})$ whenever ${\bm{x}}\sim_{R}{\bm{y}}$. We can see that the latter is true by noting that Definition 5 ensures $\Omega({\bm{x}},R)=\Omega({\bm{y}},R)$ whenever ${\bm{x}}\sim_{R}{\bm{y}}$ so that $B_{R}^{*}({\bm{x}})=B_{R}^{*}({\bm{y}})$ by Definition 17. To show the former, note that by Definition 5, ${\bm{x}}\lesssim_{R}{\bm{y}}\Rightarrow\Omega({\bm{y}},R)\subseteq\Omega({\bm{x}},R)$, which by Definition 17 implies that $B_{R}^{*}({\bm{x}})\leq B_{R}^{*}({\bm{y}})$. ∎
Theorem 2.
For every $\alpha\in[0,1)$ and total preorder $R$, $B_{R}^{*}$ is the conditionally optimal bound for $R$ provided that $\mathcal{F}(\Omega({\bm{x}},R),\alpha)$ is non-empty.
Proof.
We begin by noting that, according to Lemma 4, $B_{R}^{*}$ is consistent with $R$. Thus, by Definition 16, it suffices to show that for any valid bound $B_{R}$, consistent with $R$, and $\forall{\bm{x}}\in\Omega,B^{*}_{R}({\bm{x}})\geq B_{R}({\bm{x}})$. This would imply that $B_{R}$ is either strictly worse than $B_{R}^{*}$ for some ${\bm{x}}\in\Omega$, or that $B_{R}=B_{R}^{*}$. To that end, suppose to the contrary that there exists some valid $B_{R}$, also consistent with $R$, such that $\exists{\bm{x}}\in\Omega:B_{R}^{*}({\bm{x}})<B_{R}({\bm{x}})$. Now let $T$ be any total order agreeing with $R$ and take ${\bm{y}}\in\Omega$ such that ${\bm{y}}\sim_{R}{\bm{x}}$ and $\Omega({\bm{y}},T)=\Omega({\bm{y}},R)=\Omega({\bm{x}},R)$. We are guaranteed that such a $T$ exists by Propositions 5, and we are guaranteed that such a $\bm{y}$} exists by Proposition 7. Since $T$ agrees with $R$ we know by Definition 15 that $B_{R}$ is consistent with $T$. Notice also that Theorem 1 ensures that $B_{T}^{*}$ is conditionally optimal for $T$.
Now since ${\bm{x}}\sim_{R}{\bm{y}}$ and $B_{R}^{*}({\bm{x}})<B_{R}({\bm{x}})$, it’s clear by Lemma 3 that
$B_{R}^{*}({\bm{y}})=B_{R}^{*}({\bm{x}})<B_{R}({\bm{x}})=B_{R}({\bm{y}}).$ (19)
And because $\Omega({\bm{y}},T)=\Omega({\bm{y}},R)$ we also have by Definition 17 that $B_{T}^{*}({\bm{y}})=B_{R}^{*}({\bm{y}})$. So it must be the case that $B_{T}^{*}({\bm{y}})<B_{R}({\bm{y}})$ as well, which means that there is a bound consistent with $T$ (the bound $B_{R}$) that for sample $\bm{y}$} improves on the bound given by $B_{T}^{*}$. But this contradicts the fact that $B_{T}^{*}$ is conditionally optimal for order $T$. Thus, it cannot be the case that $B_{R}^{*}({\bm{x}})<B_{R}({\bm{x}})$. ∎
4 Pointwise-Optimality
The theory of Learned-Miller [2] establishes that there can exist no bound that is optimal for every sample in $\Omega$. Nevertheless, it remains possible to derive the highest possible bound value that can be assigned to any given sample by any valid bound.
Definition 18.
The optimal bound for sample ${\bm{x}}\in\Omega$, denoted $B^{*}({\bm{x}})$, is the highest bound value assigned to $\bm{x}$} by any valid bound.
Lemma 5.
For any ${\bm{x}}\in\Omega$,
$B^{*}({\bm{x}})=\min\{E[F]:F\in\mathcal{F}(\{{\bm{x}}\},\alpha)\}.$ (20)
Proof.
From Definition 18 and Theorem 1, it will suffice to find the highest bound value for $\bm{x}$} among all pessimal bounds (see Definition 17). These bounds amount to finding the distribution $F^{*}$ achieving minimum mean among all distributions in a set of the form $\{F:F\in\mathcal{F}(\Omega({\bm{x}},T),\alpha)\}$ for some total order $T$. The conclusion follows by noticing that each of these sets contains the set $\{F:F\in\mathcal{F}(\{{\bm{x}}\},\alpha)\}$, which corresponds to any total order that places $\bm{x}$} last in its ordering. ∎
Theorem 3.
The optimal bound for homogeneous $\bm{S}_{i}\in\Omega$ is given by
$B^{*}(\bm{S}_{i})=S_{\min}(1-\sqrt[n]{\alpha})+S_{i}\sqrt[n]{\alpha}.$ (21)
Proof.
From Lemma 5, we know that $B^{*}(\bm{S}_{i})=\min\{E[F]:F\in\mathcal{F}(\{\bm{S}_{i}\},\alpha)\}$. Let $p_{0}=\sqrt[n]{\alpha}$. Every distribution $G_{p}\in\mathcal{G}(\{\bm{S}_{i}\},\alpha)$ places mass $p=p_{0}+\epsilon$ on support point $S_{i}$, for some $\epsilon>0$. Thus, according to Lemma 2, $B^{*}(\bm{S}_{i})$ will be equal to the limiting mean of some sequence $\{G_{p_{0}+\epsilon}\}_{\epsilon\downarrow 0}$. Let $H_{p}\in\mathcal{G}(\{\bm{S}_{i}\},\alpha)$ be the distribution that places mass $p$ at support point $S_{i}$, and places the remaining mass at $S_{\min}$. It is clear that 1) $G_{p}\in\mathcal{G}(\{\bm{S}_{i}\},\alpha)\Rightarrow H_{p}\in\mathcal{G}(\{\bm{S}_{i}\},\alpha)$ 2) every sequence $\{G_{p_{0}+\epsilon}\}_{\epsilon\downarrow 0}$ is bounded below by the corresponding sequence $\{H_{p_{0}+\epsilon}\}_{\epsilon\downarrow 0}$, and 3) $E[H_{p}]\leq E[G_{p}]$. Therefore,
$B^{*}(\bm{S}_{i})=\lim_{\epsilon\rightarrow 0}E[H_{p_{0}+\epsilon}]=S_{\min}(1-\sqrt[n]{\alpha})+S_{i}\sqrt[n]{\alpha}.$ (22)
5 Monotone Orders
In this section we show that the low and high lexicographic orders are extremal in the sense that the conditionally-optimal bound for any monotone order evaluated at a given sample will fall between the values given for nearby samples by bounds conditionally-optimal for the low and high lexicographic orders.
Proposition 8.
The low and high lexicographic orders are monotone.
Proof.
We prove the proposition for $T_{\ell}$; the proof for $T_{h}$ is similar. Suppose that for any $\bm{x},\bm{y}\in\Omega$, $\bm{x}\leq\bm{y}$ but $\neg(\bm{x}\leq_{T_{\ell}}\bm{y})$. Since $\neg(\bm{x}\leq_{T_{\ell}}\bm{y})$ we know by definition of $T_{\ell}$ that $L_{1}(\bm{x},\bm{y})=0$, which implies that $\exists i\in[n]:x_{(i)}>y_{(i)}$, i.e. $\neg(\bm{x}\leq\bm{y})$. But this contradicts the assumption that $\bm{x}\leq\bm{y}$. Thus, we must have $\bm{x}\leq\bm{y}\Rightarrow\bm{x}\leq_{T}\bm{y}$. ∎
Proposition 9.
Let $\bm{y}\in\Omega$ be arbitrary and $\bm{S}_{i}\in\Omega$ be any homogeneous sample. For the low lexicographic order $T_{\ell}$ we have
$\bm{S}_{i}\leq_{T_{\ell}}\bm{y}\Leftrightarrow\bm{S}_{i}\leq\bm{y}.$ (23)
Proof.
Since $T_{\ell}$ is monotone, it suffices to show that $\bm{S}_{i}\leq_{T_{\ell}}\bm{y}\Rightarrow\bm{S}_{i}\leq\bm{y}$. Suppose to the contrary that $\bm{S}_{i}\leq_{T_{\ell}}\bm{y}$ but $\neg(\bm{S}_{i}\leq\bm{y})$. Since $\neg(\bm{S}_{i}\leq\bm{y})$ we know that $\exists j\in[n]:S_{i}>y_{(j)}$, which in turn implies that $\forall k\leq j,S_{i}>y_{(k)}$. But because $\bm{S}_{i}\leq_{T_{\ell}}\bm{y}$, we have that $L_{1}(\bm{S}_{i},\bm{y})=1$. This implies in particular that $S_{i}\leq y_{(1)}$, which leads to a contradiction. Thus, it must be the case that $\bm{S}_{i}\leq_{T_{\ell}}\bm{y}\Rightarrow\bm{S}_{i}\leq\bm{y}$. ∎
Proposition 10.
Let $\bm{y}\in\Omega$ be arbitrary and $\bm{S}_{i}\in\Omega$ be any homogeneous sample. For the high lexicographic order $T_{h}$ we have
$\bm{S}_{i}\leq_{T_{h}}\bm{y}\Leftrightarrow\bm{S}_{i}=\bm{y}\vee S_{i}<y_{(n)}.$ (24)
Proof.
The equivalence is clearly true when $\bm{S}_{i}=\bm{y}$, so we proceed under the assumption that $\bm{S}_{i}\neq\bm{y}$. We begin by showing that $\bm{S}_{i}\leq_{T_{h}}\bm{y}\Rightarrow S_{i}<y_{(n)}$. Suppose to the contrary that $\bm{S}_{i}\leq_{T_{h}}\bm{y}$ but $y_{(n)}\leq S_{i}$. Since $\bm{S}_{i}$ is homogeneous it must also be the case that $\bm{y}\leq\bm{S}_{i}$. Combined with the assumption that $\bm{S}_{i}\neq\bm{y}$, this implies that $\bm{y}<\bm{S}_{i}$. But since $T_{h}$ is monotone we also have that $\bm{y}<\bm{S}_{i}\Rightarrow\bm{y}<_{T_{h}}\bm{S}_{i}$, which contradicts the original assumption that $\bm{S}_{i}\leq_{T_{h}}\bm{y}$. Thus, $\bm{S}_{i}\leq_{T_{h}}\bm{y}\Rightarrow S_{i}<y_{(n)}$.
Next we show that $S_{i}<y_{(n)}\Rightarrow\bm{S}_{i}\leq_{T_{h}}\bm{y}$. From $S_{i}<y_{(n)}$ it is immediately clear that $H_{n}(\bm{S}_{i},\bm{y})=1$, which implies by definition that $\bm{S}_{i}\leq_{T_{h}}\bm{y}$. Thus, $S_{i}<y_{(n)}\Rightarrow\bm{S}_{i}\leq_{T_{h}}\bm{y}$. ∎
Theorem 4.
For any monotone total order $T$ and homogeneous sample $\bm{S}_{i}\in\Omega$ we have
$\Omega(\bm{S}_{i},T_{\ell})\subseteq\Omega(\bm{S}_{i},T)\subseteq\Omega(\bm{S}_{i},T_{h}).$ (25)
Proof.
We first show that $\Omega(\bm{S}_{i},T_{\ell})\subseteq\Omega(\bm{S}_{i},T)$. Assume to the contrary that $\exists\bm{y}\in\Omega$ such that $\bm{y}\in\Omega(\bm{S}_{i},T_{\ell})$ and $\bm{y}\not\in\Omega(\bm{S}_{i},T)$, the latter of which implies $\neg(\bm{S}_{i}\leq_{T}\bm{y})$. Since $T$ is monotone we have $\neg(\bm{S}_{i}\leq_{T}\bm{y})\Rightarrow\neg(\bm{S}_{i}\leq\bm{y})$. Therefore, it must be the case that $\neg(\bm{S}_{i}\leq\bm{y})$. On the other hand, since $\bm{y}\in\Omega(\bm{S}_{i},T_{\ell})$ we know by Proposition 9 that $\bm{S}_{i}\leq\bm{y}$, which leads to a contradiction. Thus, $\Omega(\bm{S}_{i},T_{\ell})\subseteq\Omega(\bm{S}_{i},T)$.
Next, we show that $\Omega(\bm{S}_{i},T)\subseteq\Omega(\bm{S}_{i},T_{h})$. Suppose to the contrary that $\exists\bm{y}\in\Omega:\bm{y}\in\Omega(\bm{S}_{i},T)$ and $\bm{y}\not\in\Omega(\bm{S}_{i},T_{h})$. Since $\bm{y}\not\in\Omega(\bm{S}_{i},T_{h})$, we know by Proposition 10 that $\bm{y}\neq\bm{S}_{i}$ and $y_{(n)}\leq S_{i}$, which together imply that $\bm{y}<\bm{S}_{i}$. And since $T$ is monotone, $\bm{y}<\bm{S}_{i}$ implies that $\bm{y}<_{T}\bm{S}_{i}$. On the other hand, $\bm{y}\in\Omega(\bm{S}_{i},T)$ implies that $\bm{S}_{i}\leq_{T}\bm{y}$, which leads to a contradiction. Thus, we have shown that $\Omega(\bm{S}_{i},T)\subseteq\Omega(\bm{S}_{i},T_{h})$. ∎
Theorem 5.
Let $T$ be any monotone total order, $B_{T}^{*}$ the conditionally optimal bound for $T$, and $\bm{x}\in\Omega$ an arbitrary sample. If $S_{i}\in S$ is such that $\bm{S}_{i}\leq_{T}\bm{x}\leq_{T}\bm{S}_{i+1}$, then
$B_{T_{h}}^{*}(\bm{S}_{i})\leq B_{T}^{*}(\bm{x})\leq B_{T_{\ell}}^{*}(\bm{S}_{i+1}).$ (26)
Proof.
We begin by stating some facts
  1. 1.
    For upper sets $\Omega_{1}\subseteq\Omega_{2}\subseteq\Omega$ we have $\mathcal{F}(\Omega_{1})\subseteq\mathcal{F}(\Omega_{2})$: Take any $F\in\mathcal{F}(\Omega_{1})$. By construction, $P_{F}[\Omega_{1}]>\alpha$. But since $\Omega_{1}\subseteq\Omega_{2}$, it must also be the case that $P_{F}[\Omega_{2}]>\alpha$, which implies that $F\in\mathcal{F}(\Omega_{2})$ as well.
  2. 2.
    Since $B_{T}^{*}$ is consistent with $T$, and $\bm{S}_{i}\leq_{T}\bm{x}\leq_{T}\bm{S}_{i+1}$, we know that $B_{T}^{*}(\bm{S}_{i})\leq B_{T}^{*}(\bm{x})\leq B_{T}^{*}(\bm{S}_{i+1})$.
  3. 3.
    For any subsets of distributions $\mathcal{F}_{1}\subseteq\mathcal{F}_{2}\subseteq\mathcal{F}$ we have that that $\min\{E[F]:F\in\mathcal{F}_{2}\}\leq\min\{E[F]:F\in\mathcal{F}_{1}\}$.
  4. 4.
    For any order $T^{\prime}$ we know by Theorem 1 that $B_{T^{\prime}}^{*}(\bm{S}_{i})=\min\{E[F]:F\in\mathcal{F}(\Omega(\bm{S}_{i},T^{\prime}),\alpha)\}$.
We now show that $B_{T}^{*}(\bm{x})\leq B_{T_{\ell}}^{*}(\bm{S}_{i+1})$. By Theorem 4 we have that $\Omega(\bm{S}_{i},T_{\ell})\subseteq\Omega(\bm{S}_{i},T)$. Thus, according to (1), $\mathcal{F}(\Omega(\bm{S}_{i},T_{\ell}),\alpha)\subseteq\mathcal{F}(\Omega(\bm{S}_{i},T),\alpha)$. Finally, from (2-4) we conclude that
$\begin{array}[]{rcl}B_{T}^{*}(\bm{x})&\leq&B_{T}^{*}(\bm{S}_{i+1})\\ &=&\min\{E[F]:F\in\mathcal{F}(\Omega(\bm{S}_{i+1},T),\alpha)\}\\ &\leq&\min\{E[F]:F\in\mathcal{F}(\Omega(\bm{S}_{i+1},T_{\ell}),\alpha)\}\\ &=&B_{T_{\ell}}^{*}(\bm{S}_{i+1})\\ \end{array}.$ (27)
We take a similar approach to show that $B_{T_{h}}^{*}(\bm{S}_{i})\leq B_{T}^{*}(\bm{x})$. Again by Theorem 4 we know that $\Omega(\bm{S}_{i},T)\subseteq\Omega(\bm{S}_{i},T_{h})$, which combined with (1) implies that $\mathcal{F}(\Omega(\bm{S}_{i},T),\alpha)\subseteq\mathcal{F}(\Omega(\bm{S}_{i},T_{h}),\alpha)$. Invoking (2-4) gives
$\begin{array}[]{rcl}B_{T_{h}}^{*}(\bm{S}_{i})&=&\min\{E[F]:F\in\mathcal{F}(\Omega(\bm{S}_{i},T_{h}),\alpha)\}\\ &\leq&\min\{E[F]:F\in\mathcal{F}(\Omega(\bm{S}_{i},T),\alpha)\}\\ &=&B_{T}^{*}(\bm{S}_{i})\\ &\leq&B_{T}^{*}(\bm{x})\\ \end{array}.$ (28)
Corollary 1.
For any homogeneous sample $S_{i}\in S$, $B_{T_{h}}^{*}(\bm{S}_{i})$ is the weakest (lowest) and $B_{T_{\ell}}^{*}(\bm{S}_{i})$ the strongest (highest) among all bounds conditionally optimal with respect to a monotone total order.
Proof.
Let $T$ be any monotone order and $B_{T}^{*}$ the conditionally optimal bound for that order. According to Theorem 5 we have simultaneously that
$B_{T_{h}}^{*}(\bm{S}_{i-1})\leq B_{T}^{*}(\bm{S}_{i})\leq B_{T_{\ell}}^{*}(\bm{S}_{i}),$ (29)
and
$B_{T_{h}}^{*}(\bm{S}_{i})\leq B_{T}^{*}(\bm{S}_{i})\leq B_{T_{\ell}}^{*}(\bm{S}_{i+1}),$ (30)
which implies that
$B_{T_{h}}^{*}(\bm{S}_{i})\leq B_{T}^{*}(\bm{S}_{i})\leq B_{T_{\ell}}^{*}(\bm{S}_{i}).$ (31)
5.1 Calculating bounds for homogeneous samples
Lemma 6.
Bound $B_{T_{\ell}}^{*}$ is conditionally optimal at the $1-\alpha$ level with respect to $\mathcal{F}$ only if $\forall S_{i}\in S$,
$B_{T_{\ell}}^{*}(\bm{S}_{i})=S_{\min}(1-\sqrt[n]{\alpha})+S_{i}\sqrt[n]{\alpha}.$ (32)
Proof.
Fix $\bm{S}_{i}$. Learned-Miller [2] Lemma 2.2 establishes that the result is true for $S_{i}=S_{\min}$, so we assume that $S_{i}\neq S_{\min}$. By Theorem 1 it will suffice to show that $\min\{E[F]:F\in\mathcal{F}(\Omega(\bm{S}_{i},T_{\ell}),\alpha)\}=\mu$ where $\mu=S_{\min}(1-\sqrt[n]{\alpha})+S_{i}\sqrt[n]{\alpha}$.
On one hand we have by Proposition 9 that for all $\bm{y}\in\Omega$, $\bm{S}_{i}\leq_{T_{\ell}}\bm{y}\Rightarrow\bm{S}_{i}\leq\bm{y}$, i.e. all samples in $\Omega(\bm{S}_{i},T_{\ell})$ comprise support values greater than or equal to $S_{i}$. Therefore, for every $G\in\mathcal{G}(\Omega(\bm{S}_{i},T_{\ell}),\alpha),E[G]>\mu$ since $G$ must assign mass exceeding $\alpha$ to samples in $\Omega(\bm{S}_{i},T_{\ell})$, which implies that the mass assigned to support values $S_{i}$ or higher must exceed $\sqrt[n]{\alpha}$. Thus, every sequence of distributions from $\mathcal{G}(\Omega(\bm{S}_{i},T_{\ell}),\alpha)$ must converge in mean to some value no less than $\mu$ so that, by Proposition 2, $\min\{E[F]:F\in\mathcal{F}(\Omega(\bm{S}_{i},T_{\ell}),\alpha)\}\geq\mu$. On the other hand, there exists a $j\in\mathbb{N}$ for which we can construct a sequence of distributions $G_{j},G_{j+1}\ldots\in\mathcal{G}(\Omega(\bm{S}_{i},T_{\ell}),\alpha)$ such that for each $i\geq j$, $G_{i}(S_{\min})=(1-\sqrt[n]{\alpha})-1/i$ and $G_{i}(S_{i})=\sqrt[n]{\alpha}+1/i$. In this case we have that $\forall\delta>0,\exists N\in\mathbb{N}:\forall n>N,E[G_{n}]-\mu<\delta$, which means that the sequence converges in mean to $\mu$. Again by Proposition 2, this implies that there exists a distribution in $\mathcal{F}(\Omega(\bm{S}_{i},T_{\ell}),\alpha)$ with mean $\mu$ so that $\min\{E[F]:F\in\mathcal{F}(\Omega(\bm{S}_{i},T_{\ell}),\alpha)\}\leq\mu$. ∎
Lemma 7.
Bound $B_{T_{h}}^{*}$ is conditionally optimal at the $1-\alpha$ level with respect to $\mathcal{F}$ only if $\forall S_{i}\in S$, $i\neq 0$,
$S_{\min}\sqrt[n]{1-\alpha}+S_{i}(1-\sqrt[n]{1-\alpha})\leq B_{T_{h}}^{*}(\bm{S}_{i})\leq S_{\min}\sqrt[n]{1-\alpha}+S_{i+1}(1-\sqrt[n]{1-\alpha}).$ (33)
Proof.
From Learned-Miller [2] Lemma 2.2, the result is true for $S_{i}=S_{\min}$, so we fix $\bm{S}_{i}\neq\bm{S}_{\min}$. By Theorem 1 we need only show that $\min\{E[F]:F\in\mathcal{F}(\Omega(\bm{S}_{i},T_{h}),\alpha)\}=\mu$ falls within the interval defined by Inequality 33. To that end, by Proposition 9, it will suffice to show the following two properties. On one hand, $\forall H\in\mathcal{G}(\Omega(\bm{S}_{i},T_{h}),\alpha)\}$, $E[H]>S_{\min}\sqrt[n]{1-\alpha}+S_{i}(1-\sqrt[n]{1-\alpha})$. And on the other hand, there exists a sequence of distributions in $\mathcal{G}(\Omega(\bm{S}_{i},T_{h}),\alpha)$ whose limiting mean is bounded above by $S_{\min}\sqrt[n]{1-\alpha}+S_{i+1}(1-\sqrt[n]{1-\alpha})$.
Fix $S_{i}\in S$, $i>0$, and let $H\in\mathcal{G}(\Omega(\bm{S}_{i},T_{h}),\alpha)$ be arbitrary. Define transformation $\phi:S\rightarrow S$ such that for each $S_{k}\in S$, $\phi(S_{k})=S_{\min}$ if $S_{k}<S_{i}$ and $\phi(S_{k})=S_{i+1}$ if $S_{k}\geq S_{i}$. Furthermore, define $G_{H}\in\mathcal{F}$ to be the distribution that, for each $S_{k}\in S$, transfers all mass from $S_{k}$ to $\phi(S_{k})$. Note that $\forall A\in\Omega,P_{G_{H}}[\phi(A)]\geq P_{H}[A]$. Let $A=\Omega(\bm{S}_{i},T_{h})$. We next show that $P_{G_{H}}[A]>\alpha$ so that $G_{H}\in\mathcal{G}(\Omega(\bm{S}_{i},T_{h}),\alpha)$. To that end, it will suffice to show that $\phi(A)\subseteq A$ so that $P_{G_{H}}[A]\geq P_{G_{H}}[\phi(A)]\geq P_{H}[A]>\alpha$. Take any $\bm{x}\in A$. If $\bm{x}=\bm{S}_{i}$, then $\phi(\bm{x})=\bm{S}_{i+1}\in A$. Otherwise, by Proposition 10, $x_{(n)}\geq S_{i}$ and so $\phi(x_{(n)})=S_{i+1}>S_{i}$, which also implies that $\phi(\bm{x})\in A$. Thus, $\phi(A)\subseteq A$.
Now define $G_{H}^{\prime}$ as the distribution that results from $G_{H}$ after transferring all mass in $G_{H}$ from $S_{i+1}$ to $S_{i}$. It’s clear by construction that $H$ stochastically dominates $G_{H}^{\prime}$ so that $E[G_{H}^{\prime}]\leq E[H]$. Notice that both $G_{H}$ and $G_{H}^{\prime}$ can be reparameterized by $p$, where fraction $p$ of the mass in the distribution is placed at $S_{\min}$ and the rest is placed at the other atom in their support. We say that $G_{p}$ and $G_{p}^{\prime}$ are achievable if $\exists H\in\mathcal{G}(\Omega(\bm{S}_{i},T_{h}),\alpha)$ such that $G_{H}=G_{p}$. Under this parameterization we have $E[G_{p}]=pS_{\min}+(1-p)S_{i+1}$ and $E[G_{p}^{\prime}]=pS_{\min}+(1-p)S_{i}$.
Finally, define $p^{*}=\sqrt[n]{1-\alpha}$. Notice that the only sample from $\phi(\Omega)$ not in $\Omega(\bm{S}_{i},T_{h})$ is $\bm{S}_{\min}$. Thus, the pair $G_{p}^{\prime}$ and $G_{p}$ is achievable iff we have that $p<p^{*}$. We have argued that for every $H\in\mathcal{G}(\Omega(\bm{S}_{i},T_{h}),\alpha)$ there exists an achievable $G_{p}^{\prime}$ such that $E[G_{p}^{\prime}]<E[H]$. Therefore, $\forall H\in\mathcal{G}(\Omega(\bm{S}_{i},T_{h}),\alpha),E[G_{p^{*}}^{\prime}]<E[H]$. Now define the sequence of distributions $G_{p^{*}-1/j}$ for each $j>\lceil 1/p^{*}\rceil$. By construction, each element in the sequence is achievable and the sequence converges to $G_{p^{*}}$. We have also argued that corresponding to each $G_{p^{*}-1/j}$ there exists $H\in\mathcal{G}(\Omega(\bm{S}_{i},T_{h}),\alpha)$ such that $E[H]\leq E[G_{p^{*}-1/j}]$. Therefore, there exists a sequence of distributions in $\mathcal{G}(\Omega(\bm{S}_{i},T_{h}),\alpha)$ whose means are bounded from above by the means of the sequence of distributions $G_{p^{*}-1/j}$. ∎
6 Refinements of $\mathcal{F}$
Definition 19.
For any $G,G^{\prime}\in\mathcal{F}$ and $C\subseteq S$, we say that $G$ and $G^{\prime}$ agree pointwise on $C$ if $\forall x\in C,P_{G}[X=x]=P_{G^{\prime}}[X=x]$ and cumulatively on $C$ if $\forall x\in C,P_{G}[{\bm{X}}\leq{\bm{x}}]=P_{G^{\prime}}[{\bm{X}}\leq{\bm{x}}]$.
Definition 20.
For arbitrary $C\in S$ define the refinement $\mathcal{F}_{C}=\{F\in\mathcal{F}:\forall S_{i}\in S\setminus C,P_{F}[S_{i}]=0\}$. For arbitrary $\Omega^{\prime}\subseteq\Omega$, we further define refinements
$\mathcal{G}_{C}(\Omega^{\prime},\alpha)=\mathcal{F}_{C}\cap\mathcal{G}(\Omega^{\prime},\alpha),$ (34)
and
$\mathcal{F}_{C}(\Omega^{\prime},\alpha)=\texttt{cl}(\mathcal{G}_{C}(\Omega^{\prime},\alpha)).$ (35)
Definition 21.
For each $C\subseteq S$, we denote by $C^{+}$ the augmentation
$C^{+}=C\cup\{S_{\min}\}\cup\{S_{i+1}:S_{i}\in C,S_{i}\neq S_{\max}\}.$ (36)
Lemma 8.
Fix ${\bm{x}}\in\Omega$, $\alpha\in[0,1)$, and total preorder $R$. Suppose that there exists a set $C\subseteq S$ such that for any $G,H\in\mathcal{F}$ agreeing cumulatively and pointwise on $C$, $P_{G}[\Omega({\bm{x}},R)]=P_{H}[\Omega({\bm{x}},R)]$. Then we have that
$B_{R}^{*}({\bm{x}})=\min\{E[F]:F\in\mathcal{F}_{C^{+}}(\Omega(\bm{x},R),\alpha)\}.$ (37)
Proof.
From Theorem 2 and Lemma 2 we have:
  1. 1.
    For every sequence $\{G_{i}\}_{i>0}$, $G_{i}\in\mathcal{G}(\Omega({\bm{x}},R),\alpha)$, $\lim_{i\rightarrow\infty}E[G_{i}]\geq B_{R}^{*}({\bm{x}})$.
  2. 2.
    There exists a sequence of distributions $\{G_{i}^{*}\}_{i>0}$, $G_{i}^{*}\in\mathcal{G}(\Omega({\bm{x}},R),\alpha)$, such that $\lim_{i\rightarrow\infty}E[G_{i}^{*}]=B_{R}^{*}({\bm{x}})$
We will first show that for every $G\in\mathcal{G}(\Omega({\bm{x}},R),\alpha)$, there exists an $H\in\mathcal{G}_{C^{+}}(\Omega({\bm{x}},R),\alpha)$ such that $E[H]\leq E[G]$. This establishes that, for the sequence $\{G_{i}^{*}\}_{i>0}$ in particular, there exists corresponding sequence $\{H_{i}^{*}\}_{i>0}$, $H_{i}^{*}\in\mathcal{G}_{C^{+}}(\Omega({\bm{x}},R),\alpha)$, such that $\lim_{i\rightarrow\infty}E[H_{i}^{*}]\leq\lim_{i\rightarrow\infty}E[G_{i}^{*}]$, which would imply by Property 2 that $\lim_{i\rightarrow\infty}E[H_{i}^{*}]\leq B_{R}^{*}({\bm{x}})$. On the other hand, since $\mathcal{G}_{C^{+}}(\Omega({\bm{x}},R),\alpha)\subseteq\mathcal{G}(\Omega({\bm{x}},R),\alpha)$, Property 1 establishes that $\lim_{i\rightarrow\infty}E[H_{i}^{*}]\geq B_{R}^{*}({\bm{x}})$. Thus, it will follow that the distribution with lowest mean in $\mathcal{F}(\Omega(\bm{x},R),\alpha)$ is also in $\mathcal{F}_{C^{+}}(\Omega(\bm{x},R),\alpha)$.
Fix $G\in\mathcal{G}(\Omega({\bm{x}},R),\alpha)$ and construct distribution $H$ as follows. Let $k=|C|$, $s_{j}$ denote the $j$th largest element in $C$, and $s_{j}^{\prime}$ be the smallest element in $S$ such that $s_{j}<s_{j}^{\prime}$. $H$ agrees both pointwise and cumulatively with $G$ on $C$ and
  1. 1.
    If $s_{1}>S_{\min}$, then $H(S_{\min})=P_{G}[X<s_{1}]$.
  2. 2.
    If $s_{k}<S_{\max}$, then $H(s_{k}^{\prime})=1-P_{G}[X\leq s_{k}]$.
  3. 3.
    $\forall j\in\{1,\ldots,k-1\}$, if $s_{j}^{\prime}\neq s_{j+1}$, then $H(s_{j}^{\prime})=P_{G}[X<s_{j+1}]-P_{G}[X\leq s_{j}]$.
In words, Property 3 states that $H(s_{j}^{\prime})$ is equal to the mass that $G$ assigns to support points above $s_{j}$ but below $s_{j+1}$. Because $G$ and $H$ agree pointwise and cumulatively on $C$, we have by assumption that $P_{G}[\Omega({\bm{x}},R)]=P_{H}[\Omega({\bm{x}},R)]$, which implies that $H\in\mathcal{G}_{C^{+}}(\Omega(\bm{x},R),\alpha)$. And by construction we also have that $E[H]\leq E[G]$. ∎
Theorem 6.
Let $C=\{x_{(i)}\}$. For fixed ${\bm{x}}\in\Omega$ and $\alpha\in[0,1)$,
$B_{R_{i}}^{*}({\bm{x}})=\min\{E[F]:F\in\mathcal{F}_{C^{+}}(\Omega(\bm{x},R_{i}),\alpha)\}.$ (38)
Proof.
From Lemma 8 it will suffice to show that for any two distributions $G,H\in\mathcal{F}$ agreeing cumulatively and pointwise on $C$, $P_{G}[\Omega({\bm{x}},R_{i})]=P_{H}[\Omega({\bm{x}},R_{i})]$. By Definition 10 and the fact that $H$ agrees cumulatively with $G$ at $x_{(i)}\in C$ we have
$\begin{array}[]{rcl}P_{G}[\Omega({\bm{x}},R_{i})]&=&P_{G}[{\bm{x}}\lesssim_{R_{i}}{\bm{Y}}]\\ &=&P_{G}[x_{(i)}\leq Y_{(i)}]\\ &=&P_{H}[x_{(i)}\leq Y_{(i)}]\\ &=&P_{H}[{\bm{x}}\lesssim_{R_{i}}{\bm{Y}}]\\ &=&P_{H}[\Omega({\bm{x}},R_{i})]\end{array}.$ (39)
Theorem 7.
Let $c=(S_{\max}-S_{\min})/(m-1)$. For each ${\bm{x}}\in\Omega$, $\alpha\in[0,1)$, and $\epsilon>0$, there exists a polynomial-time $(c+\epsilon)$-approximation for calculating $B_{R_{i}}^{*}({\bm{x}})$.
Proof.
Fix ${\bm{x}}\in\Omega$. Let $C=\{x_{(i)}\}$, $C^{-}=\{S_{\min},x_{(i)}\}$, and, by Definition 21, we have $C^{+}=\{S_{\min},x_{(i)},x_{(i)}^{\prime}\}$, where $x_{(i)}^{\prime}$ is $S_{\max}$ if $x_{(i)}=S_{\max}$ or otherwise the least element of $S$ that is greater than $x_{(i)}$. Theorem 6 establishes that there exists a distribution $F^{*}\in\mathcal{F}_{C^{+}}(\Omega({\bm{x}},R_{i}),\alpha)$ whose mean achieves $B_{R_{i}}^{*}(\bm{x})$. And by Lemma 1, we also have that $\forall\varepsilon>0,\exists G^{*}\in\mathcal{G}_{C^{+}}(\Omega({\bm{x}},R_{i}),\alpha):|E[F^{*}]-E[G^{*}]|<\varepsilon$.
Corresponding to each $G\in\mathcal{G}_{C^{+}}(\Omega({\bm{x}},R_{i}),\alpha)$ is a distribution $H_{p}\in\mathcal{F}_{C^{-}}$ where $H_{p}$ is equal to $G$, except that all mass at $G(x_{(i)}^{\prime})$ is transferred to $x_{(i)}$ such that $P_{H_{p}}[x_{(i)}]=p$ and $P_{H_{p}}[S_{\min}]=1-p$. Notice that, by construction,
$\begin{array}[]{rcl}E[G]-E[H_{p}]&\leq&p(x_{(i)}^{\prime}-x_{(i)})\\ &\leq&(x_{(i)}^{\prime}-x_{(i)})\\ &\leq&\frac{S_{\max}-S_{\min}}{m-1}\\ \end{array}.$ (40)
Also by construction and Definition 10, $P_{G}[x_{(i)}\leq X_{(i)}]=P_{H_{p}}[x_{(i)}\leq X_{(i)}]$. Therefore, $H_{p}\in\mathcal{G}_{C^{+}}(\Omega({\bm{x}},R_{i}),\alpha)$. Take $H_{p^{*}}$ to be the distribution in $\mathcal{F}_{C^{-}}$ corresponding to $G^{*}$. By letting $\varepsilon\rightarrow 0$ and by Inequality 40 and the definition of $G^{*}$ we have
$|E[F^{*}]-E[H_{p^{*}}]|\leq\frac{S_{\max}-S_{\min}}{m-1},$ (41)
where $E[H_{p}]=px_{(i)}$. Thus, it remains only to approximate $p^{*}$ to within an additive factor of $\epsilon/x_{(i)}$.
According to Definition 10 we know that ${\bm{y}}\in\Omega(\bm{x},R_{i})\Leftrightarrow x_{(i)}\leq y_{(i)}$, and in order to have $x_{(i)}\leq y_{(i)}$, given support $C^{-}$, it must be the case that there are at least $i$ occurrences of $x_{(i)}$ in $\bm{y}$}, with the remaining elements of $\bm{y}$} being $S_{\min}$. Let $V_{i}$ be the event that any such $\bm{y}$} is drawn. For distribution $H_{p}$ we have, $H_{p}\in\mathcal{G}_{C^{+}}(\Omega({\bm{x}},R_{i}))\Leftrightarrow P_{H_{p}}[V_{i}]>\alpha$ where $P_{H_{p}}[V_{i}]=1-\texttt{Bin}(n-i-1;n,p)$ and Bin denotes the cumulative binomial distribution. $\texttt{Bin}(n-i-1;n,p)$ must be decreasing in $p$ since it gives the probability that the number of successes is limited to ni1n-i-1, which implies that $P_{H_{p}}[V_{i}]$ is increasing in $p$. Of course $E[H_{p}]$ is also increasing in $p$. So to approximate $p^{*}$ we seek the smallest $p$ such that $P_{H_{p}}[V_{i}]>\alpha$, i.e. $H_{p}$ remains in $\mathcal{G}_{C^{+}}(\Omega({\bm{x}},R_{i}))$. We perform binary search for $p^{*}$ on the interval $[0,1]$ as follows.
  1. 1.
    For interval $I=[a,b]$, let |I|=ba|I|=b-a, $I^{-}=[a,a+|I|/2]$, and $I^{+}=[b-|I|/2,b]$.
  2. 2.
    Initially choose $I=[0,1]$ and $\delta=\epsilon/x_{(i)}$.
  3. 3.
    While $|I|>\delta$, take $p^{\prime}=a+|I|/2$ and
    1. (a)
      If $P_{H_{p^{\prime}}}[V_{i}]<\alpha$, then set $I=I^{+}$
    2. (b)
      Otherwise set $I=I^{-}$.
  4. 4.
    Return $\min(I)$.
At every stage of the algorithm, interval $I$ contains $p^{*}$. Yet by the end of the algorithm, $|I|\leq\delta$. Thus, by choosing $p=\min(I)$, we can be sure that $|p-p^{*}|\leq\epsilon/x_{(i)}$.
The algorithm runs for $O(\log(x_{(i)}/\epsilon))$ steps. And in each step, $P_{H_{p^{\prime}}}[V_{i}]$ can be calculated to within machine precision in time polynomial in $n$ and $i$ using common numerical procedures. Thus, the overall algorithm is polynomial in the inputs $n$, $i$, $x_{(i)}$, and $\epsilon^{-1}$. ∎
Theorem 8.
Let $C=\{S_{j}\in S:\exists k\in[n]:x_{k}=S_{j}\}$. For fixed ${\bm{x}}\in\Omega$,
$B_{T_{\ell}}^{*}({\bm{x}})=\min\{E[F]:F\in\mathcal{F}_{C^{+}}(\Omega(\bm{x},T_{\ell}),\alpha)\}.$ (42)
Proof.
From Lemma 8 it will suffice to show that for any two distributions $G,H\in\mathcal{F}$ agreeing cumulatively and pointwise on $C$, $P_{G}[\Omega({\bm{x}},T_{\ell})]=P_{H}[\Omega({\bm{x}},T_{\ell})]$. According to Definition 8, we have that $G\in\mathcal{G}(\Omega(\bm{x},T_{\ell}),\alpha)\Leftrightarrow P_{G}[L_{1}({\bm{x}},{\bm{Y}})]>\alpha$. But, for each $j\in[n]$ and any ${\bm{x}},{\bm{y}}\in\Omega$, the events $(x_{(j)}<y_{(j)})$ and $(x_{(j)}=y_{(j)})$ are mutually exclusive, so from Equation 6 we have that for all $j\in[n]$
$\begin{array}[]{l}P_{G}[L_{j}({\bm{x}},{\bm{Y}})]=\\ P_{G}[x_{(j)}<Y_{(j)}]+P_{G}\left[L_{j+1}({\bm{x}},{\bm{Y}})~|~x_{(j)}=Y_{(j)}\right]P_{G}[x_{(j)}=Y_{(j)}]\end{array}.$ (43)
Therefore, applying Lemma 9, it follows that $P_{G}[L_{1}({\bm{x}},{\bm{Y}})]=P_{H}[L_{1}({\bm{x}},{\bm{Y}})]$ so that
$\begin{array}[]{rcl}P_{G}[\Omega({\bm{x}},T_{\ell})]&=&P_{G}[{\bm{x}}\leq_{T_{\ell}}{\bm{Y}}]\\ &=&P_{G}[L_{1}({\bm{x}},{\bm{Y}})]\\ &=&P_{H}[L_{1}({\bm{x}},{\bm{Y}})]\\ &=&P_{H}[{\bm{x}}\leq_{T_{\ell}}{\bm{Y}}]\\ &=&P_{H}[\Omega({\bm{x}},T_{\ell})]\end{array}.$ (44)
Appendix A Supporting Probabilistic Results
In this section we introduce some additional vector notation to ease exposition. Let ${\bm{x}}_{(:i)}$ and ${\bm{x}}_{(i:)}$ denote, respectively, the order statistics of $\bm{x}$} before and after the $i$th order statistic (inclusive), with ${\bm{x}}_{(1:)}={\bm{x}}_{(:n)}$ denoting the entire vector $\bm{x}$} sorted in increasing order.
Lemma 9.
Let ${\bm{x}}\in\Omega$ be fixed and suppose that $G,G^{\prime}\in\mathcal{F}$ agree both cumulatively and pointwise on $S_{\bm{x}}$. Then $\forall i\in[n]$ we have
$P_{G}[x_{(i)}=Y_{(i)}|{\bm{x}}_{(:i-1)}={\bm{Y}}_{(:i-1)}]=P_{G^{\prime}}[x_{(i)}=Y_{(i)}|{\bm{x}}_{(:i-1)}={\bm{Y}}_{(:i-1)}],$ (45)
and
$P_{G}[x_{(i)}<Y_{(i)}|{\bm{x}}_{(:i-1)}={\bm{Y}}_{(:i-1)}]=P_{G^{\prime}}[x_{(i)}<Y_{(i)}|{\bm{x}}_{(:i-1)}={\bm{Y}}_{(:i-1)}],$ (46)
where for each $H\in\mathcal{F}$, $P_{H}[x_{(1)}=Y_{(1)}|{\bm{x}}_{(:0)}={\bm{Y}}_{(:0)}]\equiv P_{H}[x_{(1)}=Y_{(1)}]$ and $P_{H}[x_{(1)}<Y_{(1)}|{\bm{x}}_{(:0)}={\bm{Y}}_{(:0)}]\equiv P_{H}[x_{(1)}<Y_{(1)}]$.
Proof.
First note that since $G$ and $G^{\prime}$ agree cumulatively and pointwise on $S_{\bm{x}}$ it is also the case that $\forall x\in S_{\bm{x}}$, $P_{G}[X<x]=P_{G^{\prime}}[X<x]$. It can be shown [1][Thm. 3.3.1] that for fixed $i\in[n]$ and any $H\in\mathcal{F}$, $P_{H}[{\bm{X}}_{(:i)}={\bm{x}}_{(:i)}]=I({\bm{x}}_{(:i)},i,n)$, where
$I({\bm{x}}_{(:i)},i,n)=C(i,n)\int\limits_{B_{H}({\bm{x}}_{(:i)},i)}D({\bm{u}},i,n)d{\bm{u}},$ (47)
$C(i,n)$ and $D({\bm{u}},i,n)$ are known but unimportant for our purposes, and
$\begin{array}[]{l}B_{H}({\bm{x}}_{(:i)},i)\equiv\\ \{(u_{1},\ldots,u_{i}):\forall j\leq i,u_{j}\leq u_{j+1},P_{H}[X<x_{(j)}]\leq u_{j}\leq P_{H}[X\leq x_{(j)}]\}\end{array}.$ (48)
This implies that $I({\bm{x}}_{(:i)},i,n)$ depends on $H$ only through $B_{H}({\bm{x}}_{(:i)},i,n)$, and only at the points in $S_{{\bm{x}}(:i)}$. Thus, it is clear that $B_{G}=B_{G^{\prime}}$, and therefore
$P_{G}[{\bm{X}}_{(:i)}={\bm{x}}_{(:i)}]=P_{G^{\prime}}[{\bm{X}}_{(:i)}={\bm{x}}_{(:i)}].$ (49)
The validity of Equation 45 follows by observing that
$\begin{array}[]{rcl}P_{G}[x_{(i)}=Y_{(i)}|{\bm{x}}_{(:i-1)}={\bm{Y}}_{(:i-1)}]&=&\frac{P_{G}[{\bm{x}}_{(:i)}={\bm{Y}}_{(:i)}]}{P_{G}[{\bm{x}}_{(:i-1)}={\bm{Y}}_{(:i-1)}]}\\ &=&\frac{P_{G^{\prime}}[{\bm{x}}_{(:i)}={\bm{Y}}_{(:i)}]}{P_{G^{\prime}}[{\bm{x}}_{(:i-1)}={\bm{Y}}_{(:i-1)}]}\\ &=&P_{G^{\prime}}[x_{(i)}=Y_{(i)}|{\bm{x}}_{(:i-1)}={\bm{Y}}_{(:i-1)}]\end{array}$ (50)
Since $\bm{x}$} can be arbitrary, it also follows from Equation 49 that
$\begin{array}[]{rcl}P_{G}[x_{(1)}<Y_{(1)}]&=&\sum\limits_{y\in S,y>x_{(1)}}P_{G}[Y_{(1)}=y]\\ &=&\sum\limits_{y\in S,y>x_{(1)}}P_{G^{\prime}}[Y_{(1)}=y]\\ &=&P_{G^{\prime}}[x_{(1)}<Y_{(1)}]\end{array}.$ (51)
Now for fixed $i\in\{2,\ldots,n\}$ and $H\in\mathcal{F}$ define $S^{\prime}\subseteq S$ such that $y\in S^{\prime}$ iff $x_{(i)}<y\leq S_{\max}$. Notice that, by construction,
$\bigcup_{y\in S^{\prime}}\{t:P_{H}[X<y]\leq t\leq P_{H}[X\leq y]\}=\{t:P_{H}[X<x_{(i)}]<t\leq 1\}.$ (52)
Define
$I_{H}^{\prime}(t,{\bm{x}}_{(:i-1)},i,n)=C(i,n)\int\limits_{B_{H}({\bm{x}}_{(:i-1)},i-1)}D({\bm{u}}_{(:i-1)}\oplus\{t\},i,n)d{\bm{u}}_{(:i-1)},$ (53)
where operator $\oplus$ indicates concatenation, and again the only dependence on H is through $B_{H}$. We have
$\begin{array}[]{rcl}P_{H}[X_{(i)}>x_{(i)},{\bm{X}}_{(:i-1)}={\bm{x}}_{(:i-1)}]&=&\sum\limits_{y\in S^{\prime}}I_{H}({\bm{x}}_{(:i-1)}\oplus\{y\},i,n)\\ &=&\sum\limits_{y\in S^{\prime}}\int\limits_{B_{H}^{\prime}(y)}I_{H}^{\prime}(t,{\bm{x}}_{(:i-1)},i,n)dt\\ &=&\int\limits_{B_{H}^{\prime\prime}(x_{(i)},t)}I_{H}^{\prime}(t,{\bm{x}}_{(:i-1)},i,n)dt\\ \end{array},$ (54)
where
$B_{H}^{\prime}(y)=\{t:P_{H}[X<y]\leq t\leq P_{H}[X\leq y]\},$ (55)
and
$B_{H}^{\prime\prime}(x_{(i)},t)=\{t:P_{H}[X\leq x_{(i)}]<t\leq 1\}.$ (56)
Ultimately the only dependence on $H$ in Equation 54 is through $B_{H}$ and $B_{H}^{\prime\prime}$. And, because $G$ and $G^{\prime}$ agree cumulatively on $S_{\bm{x}}$, it must be the case that $B_{G}^{\prime\prime}=B_{G^{\prime}}^{\prime\prime}$. Therefore, it is clear that
$P_{G}[x_{(i)}<Y_{(i)},{\bm{x}}_{(:i-1)}={\bm{Y}}_{(:i-1)}]=P_{G^{\prime}}[x_{(i)}<Y_{(i)},{\bm{x}}_{(:i-1)}={\bm{Y}}_{(:i-1)}],$ (57)
and the validity of Equation 46 follows by similar argument as in Equation 50. ∎